OpenStudy (anonymous):

What is 9x+8y+24=x+y in slope intercept form?

OpenStudy (stefrheart):

First subtract the 9x onto the other side

OpenStudy (anonymous):

24=-8x-7y

OpenStudy (stefrheart):

okay so now add the 8x onto the other side

OpenStudy (anonymous):

I already did 9x+8y+24=x+y -9x -9x 8y+24=-8x+y 24=-8x-7y

OpenStudy (stefrheart):

yes but you need to make it into slope intercept form which is y=mx+b

OpenStudy (anonymous):

8x+24=-7y

OpenStudy (stefrheart):

now you need to divide the -7

OpenStudy (anonymous):

Is the answer y=8x+17?

OpenStudy (stefrheart):

where did you get that

OpenStudy (anonymous):

8x+24=-7y -- -- -7 -7 -7+8x+24=y

OpenStudy (anonymous):

Subtract 24-7=17

OpenStudy (stefrheart):

it should be y=-8/7x+24

OpenStudy (anonymous):

Thank you!

OpenStudy (stefrheart):

welcome! :)